# Knowledge Sharing in Manufacturing using Large Language Models: User Evaluation and Model Benchmarking

### 10 Jan 2024 v1

@@ 2.0.3. Query Construction
3. Model Benchmarking
@@
In our benchmarking experiment, we evaluated various commercial and open-source LLMs, including OpenAI’s ChatGPT (GPT-3.5 and GPT-4 from July 20th 2023), Guanaco 65B and 35B variants (Dettmers et al., 2023) based on Meta’s Llama (Large Language Model Meta AI) (Touvron et al., 2023a), Llama2 (Touvron et al., 2023b) with 70 billion parameters, and one of its derivatives, StableBeluga2 (Mahan et al., 2023).
GPT-3, or Generative Pre-trained Transformer 3, is an autoregressive language model developed by OpenAI. This third-generation model utilizes deep learning to generate human-like text from a given prompt, capable of producing various types of data such as word lists or code lines (Floridi and Chiriatti, 2020). OpenAI recently employed a Reinforcement Learning from Human Feedback (RLHF) strategy to enhance GPT-3’s reliability, usefulness, and alignment (OpenAI, 2023). This technique relies on human labellers providing desired model behavior examples and ranking responses from the GPT-3 model to specific prompts (OpenAI, 2023). Using this feedback, OpenAI fine-tunes the GPT-3 model, creating ChatGPT, a leading language model, introduced as a prototype on November 30, 2022 (OpenAI, 2023).
On the other hand, LLaMA, launched by Meta AI in February 2023, is a contrasting open-source language model (Touvron et al., 2023a). Meta AI trained a spectrum of models with parameters ranging from 7 billion to 65 billion. Like GPT-3, LLaMA also benefited from RLHF to enhance its capabilities. The LLaMA 7B model further paved the way for Alpaca’s development, a training regimen released by Taori et al. (2023). Alpaca employs a unique ”Self-Instruct” instruction tuning method, allowing it to achieve skills similar to OpenAI’s GPT-3 series (specifically the text-davinci-003 model) at a reduced cost. More recently, (Dettmers et al., 2023) released QLoRA, an efficient finetuning approach that allows 65B parameter models to be run on a single 48Gb GPU. These models, named Guanaco, are available in the following sizes: 7B, 13B, 33B, and 65B. According to the Vicuna benchmark, the 65B and 33B parameter models reached 99.3% and 97.3% ChatGPT performance, respectively. Nevertheless, like all LLaMA derivatives, they can only be used for research, not commercial applications. However, LLaMA 2 (Touvron et al., 2023b), which was released on July 18th, 2023, is available for both commercial and research use444https://about.fb.com/news/2023/07/llama-2/—last accessed January 10, 2024.
@@
We used a web UI for LLMs555https://github.com/oobabooga/text-generation-webui/tree/main—last accessed January 10, 2024 to load and test the Guanaco model and StableBeluga2 models. The models were loaded on a pair of Nvidia A6000s with NVlink and total Video Random Access Memory (VRAM) capacity of 96 GB. The 65B model was run in 8-bit mode to fit in the available VRAM. The models were used in the chat mode, with the llama-precise preset and fixed zero seed for reproducibility. Llama2 was evaluated using the demo on huggingface666https://huggingface.co/meta-llama/Llama-2-70b-chat-hf—last accessed January 10, 2024.
@@
To rigorously assess the models, we prepared 20 questions of varying complexity based on two types of context material: formal operating material and informal issue reports. The model prompt was constructed using the above template (2.0.3). Ultimately, the difficulty of a question is a combination of the question’s complexity and the clarity of the source material. Simple questions include retrieving a single piece of information clearly stated in the context material, for example, ”At what temperature is relubrication necessary for the OKS 4220 grease?”. Conversely, difficult questions require more reasoning or comprise multiple parts, for example, ”What should I do if the central turntable is overloaded?” which has a nuanced answer dependent on several factors not clearly articulated in the context material.
@@
Every response is manually scored on factuality, completeness, and hallucinations. Factuality refers to aligning with the facts in the context material. Complete responses contain all the information relevant to the question in the context material. A hallucinated response appears grammatically and semantically coherent but is not based on the context material. The following scoring protocol is applied: one is awarded for a completely factual, complete, or hallucinated response. In contrast, a score of 0.5 is awarded for a slightly nonfactual, incomplete, or hallucinated response (e.g., the response includes four out of the five correct steps). Otherwise, a score of zero is awarded. Therefore, wrong answers are penalized heavily. If the model responds by saying it cannot answer the question and does not make any attempt to do so, it is scored zero for factuality and completeness, but no score is given for hallucination. As such, the final score for hallucination is calculated as follows:
@@
As shown in Figure 3 and Table 1, GPT-4 outperforms other models regarding factuality, completeness, and lack of hallucinations but is closely followed by StableBeluga2 and GPT-3.5. The Guanaco models, based on Llama 1, perform significantly worse. The conciseness of the responses showed a similar pattern, except that StableBeluga2 produced the shortest answers (58 words), followed closely by GPT-4 (69 words).
@@
4. User Study at the Factory
We conducted a user study on the abovementioned system we co-designed with factory employees. We recruited factory managers for the user study. The study comprised three tasks: to ask the system several questions as if they were operators, to fill in a ”yellow tag” (issue report) based on a recent issue and request a logical check, and finally, to upload new documents to the system. After each task, they were asked to provide feedback. Then, they were posed with several open questions about benefits, risks, and challenges. Finally, demographic information was requested, such as age, gender, and role.
An inductive thematic analysis of the answers to the open questions resulted in the following five themes:
Usability (5): The system is praised for its user-friendliness, speed in delivering responses, and logical control function. Users appreciate the voice command and information search capabilities (P4). However, the user interface could be improved by having attachments appear after the main text for more accessible consultation (P1).
Content (5): The system retrieves relevant information based on queries posed. Users appreciate access to relevant documents and seeing where the data is taken from. However, there’s a call for more detailed instructions to be added to the system. Also, it’s mentioned that the part of answers based on yellow tags sometimes refers to previous knowledge, causing confusion (P2).
Features (4): There are suggestions to include additional features, such as the ability to open IO, SPL, and yellow tags in .pdf format for consultation (P3). Furthermore, one participant appreciated the ability to change file names (P4). Lastly, another participant could not access the yellow tag insertion fields as the fields were in a collapsed state (P1).
Risks & Benefits (4): The system is recognized for its potential to modernize the factory and make specific tasks faster, for example, ”Greater speed in carrying out some small activities.” (P3) and ”Being able to easily find information, especially if they are in different documents.” (P4). However, concerns are raised about the safety risks if the system provides inadequate answers and the potential time-wasting if the system lacks information on specific problems (P2, P3).
Employee Acceptance & Training (2): Users hope for system acceptance among all employees. There’s also a suggestion about training and sensitization of the operators to the system’s advantages (P3). P2 implies that users may turn to a human expert if available instead of using the system as the ”[Information from the system] might be less quick, less effective, and less exhaustive than asking the expert staff present.”
5. Summary and Discussion
5.1. Model Benchmarking
@@
GPT-4 performs best across all measures but is closely followed by StableBeluga2 and GPT-3.5. Compared to GPT-4, the cost per input token for GPT-3.5 is significantly lower777https://openai.com/pricing#language-models—last accessed January 10, 2024. The higher costs of GPT-4 are partially counteracted by the response length as GPT-3.5 (and Llama 2) tended to be wordier and include additional details that were not directly requested, whereas GPT-4 (and StableBeluga2) generated more concise responses.
@@
The latest generation of open-source models based on Llama-2 demonstrates a clear jump forward relative to its predecessors based on Llama-1, which were more prone to hallucinations and exhibited poorer reasoning abilities over the context material. While open-source models like StableBeluga2 do not score as high as GPT-4, they ensure better data security, privacy, and customization. Unlike GPT-4 and GPT-3.5, they can be hosted locally and heavily customized. This can be a crucial consideration for companies with sensitive data or unique needs. It should be noted that fine-tuning GPT-3.5 is possible from August 22nd 2023, and it was announced that fine-tuning GPT-4 would be available in the fall of 2023 888https://openai.com/blog/gpt-3-5-turbo-fine-tuning-and-api-updates—last accessed January 10, 2024.
5.2. Perceptions at the Factory
@@
Users appreciate the system’s functionality and see it as a tool for modernizing factory operations and speeding up operations. They are keen on improvements to be made for better user experience and utility, especially in the areas of content, feature enhancements, and user training. However, they express concerns about potential safety risks and the efficacy of information retrieval compared to consulting expert personnel. Although the tool is seen as inferior to learning from a human expert, it can help when human experts are not present or are unfamiliar with the issue at hand.
5.3. Limitations
The participant pool for the user study comprised of factory managers and was limited in size. This potentially overlooks the varied perspectives of other stakeholders, such as factory operators. To partially mitigate this we asked the participants to pose queries as if they were operators.
@@ 5.3. Limitations
We used the same prompt for all LLMs; however, it is possible that some of the LLMs would perform better with a prompt template developed explicitly for it. Furthermore, we matched the settings of the LLMs (e.g., temperature) as closely as possible across all the tested models; however, the same settings across model types were not completely equivalent, and in the case of Llama 2, we did not have access to the presets as we did not host it locally.
@@
The study’s design did not include a comprehensive real-world evaluation involving end users in their natural working environment, as this was considered too risky for our industry partner. Such an environment might present unique challenges and considerations not addressed in this research.
Our benchmarking procedure involved only 20 questions, and a singular coder assessed the responses. This introduces the potential for bias and limits the generalizability of our findings. Furthermore, the limited number of questions may not cover the full spectrum of complexities in real-world scenarios.
These limitations suggest directions for future research, where a broader participant pool, more comprehensive prompt and model customization, and real-world evaluations could be considered. Furthermore, considering the pace at which LLMs are improving, it would be desirable to (semi)automate benchmarking.
@@ 6. Conclusion
The results demonstrated GPT-4’s superior performance over other models regarding factuality, completeness, and minimal hallucinations. Interestingly, open source models like StableBeluga2 followed close behind. The user study highlighted the system’s user-friendliness, speed, and logical functionality. However, improvements in the user interface and content specificity were suggested, along with potential new features. Benefits included modernizing factory operations and speeding up certain tasks, though concerns about safety and efficiency were raised.
@@ Acknowledgements.
- Floridi and Chiriatti (2020) Luciano Floridi and Massimo Chiriatti. 2020. GPT-3: Its nature, scope, limits, and consequences. Minds and Machines 30, 4 (2020), 681–694.
@@
- Kernan Freire et al. (2023) Samuel Kernan Freire, Mina Foosherian, Chaofan Wang, and Evangelos Niforatos. 2023. Harnessing Large Language Models for Cognitive Assistants in Factories. In Proceedings of the 5th International Conference on Conversational User Interfaces (Eindhoven, Netherlands) (CUI ’23). Association for Computing Machinery, New York, NY, USA, Article 44, 6 pages. https://doi.org/10.1145/3571884.3604313
@@
- Lahat et al. (2023) Adi Lahat, Eyal Shachar, Benjamin Avidan, Zina Shatz, Benjamin S. Glicksberg, and Eyal Klang. 2023. Evaluating the use of large language model in identifying top research questions in gastroenterology. Scientific Reports 2023 13:1 13 (3 2023), 1–6. Issue 1. https://doi.org/10.1038/s41598-023-31412-2
@@
- Maddikunta et al. (2022) Praveen Kumar Reddy Maddikunta, Quoc-Viet Pham, Prabadevi B, N Deepa, Kapal Dev, Thippa Reddy Gadekallu, Rukhsana Ruby, and Madhusanka Liyanage. 2022. Industry 5.0: A survey on enabling technologies and potential applications. Journal of Industrial Information Integration 26 (2022), 100257. https://doi.org/10.1016/j.jii.2021.100257
@@
- OpenAI (2023) OpenAI. 2023. ChatGPT: Optimizing Language Models for Dialogue. https://openai.com/blog/chatgpt/.
@@
- Taori et al. (2023) Rohan Taori, Ishaan Gulrajani, Tianyi Zhang, Yann Dubois, Xuechen Li, Carlos Guestrin, Percy Liang, and Tatsunori B. Hashimoto. 2023. Stanford Alpaca: An Instruction-following LLaMA model. https://github.com/tatsu-lab/stanford_alpaca.
@@
- Wang et al. (2023) Xingzhi Wang, Nabil Anwer, Yun Dai, and Ang Liu. 2023. ChatGPT for design, manufacturing, and education. Procedia CIRP 119 (2023), 7–14. https://doi.org/10.1016/j.procir.2023.04.001 The 33rd CIRP Design Conference.
@@
- Wei et al. (2022) Jason Wei, Yi Tay, Rishi Bommasani, Colin Raffel, Barret Zoph, Sebastian Borgeaud, Dani Yogatama, Maarten Bosma, Denny Zhou, Donald Metzler, Ed H. Chi, Tatsunori Hashimoto, Oriol Vinyals, Percy Liang, Jeff Dean, and William Fedus. 2022. Emergent Abilities of Large Language Models. arXiv:2206.07682 [cs.CL]
@@
- Xie et al. (2023) Qianqian Xie, Weiguang Han, Xiao Zhang, Yanzhao Lai, Min Peng, Alejandro Lopez-Lira, and Jimin Huang. 2023. PIXIU: A Large Language Model, Instruction Data and Evaluation Benchmark for Finance. (6 2023). https://arxiv.org/abs/2306.05443v1
@@
- Xu et al. (2021) Xun Xu, Yuqian Lu, Birgit Vogel-Heuser, and Lihui Wang. 2021. Industry 4.0 and Industry 5.0—Inception, conception and perception. Journal of Manufacturing Systems 61 (2021), 530–535. https://doi.org/10.1016/j.jmsy.2021.10.006
- Zhang et al. (2023) Jianzhang Zhang, Yiyang Chen, Nan Niu, and Chuang Liu. 2023. A Preliminary Evaluation of ChatGPT in Requirements Information Retrieval. (4 2023). https://arxiv.org/abs/2304.12562v1

### 26 Feb 2024 v2

@@ 3.3. Query Construction
4. User Study in the Field
We conducted a user study on the system to uncover perceived benefits, usability issues, risks, and barriers to adoption. The study comprised three tasks: to ask the system several questions as if they were operators, to fill in a “yellow tag” (issue analysis report) based on a recent issue and request a logical check, and finally, to upload new documents to the system. After each task, they were asked to provide feedback. Then, after completing all tasks, the participants were posed several open questions about the system’s benefits, risks, and barriers to adoption. Finally, demographic information, such as age, gender, and role, was collected.
4.1. Participants
We recruited participants from a detergent factory, of which were managers (P1-4), and were operators (P5-9). Of the nine participants, were women, and were men. Participant age was distributed over three brackets, namely were 30–39, were 40–49, and were 50–59.
4.2. Qualitative Analysis
An inductive thematic analysis (Guest et al., 2011) of the answers to the open questions resulted in six themes discussed below.
Usability The theme of usability underlines the system’s ease of use and the need for clear instructions. Users mentioned the necessity for a “user-friendly” (P2) interface and highlighted the importance of having “more instructions and more details need to be loaded” (P1) to avoid confusion. This indicates a desire for intuitive navigation that could enable workers to use the system effectively without extensive training or referencing external help. The feedback suggests that the system already works well, as reflected in statements like “Easy-to-use system” (P3) and the system “works well” (P7).
Access to information Users appreciated the “ease of having information at hand” (P1), facilitating immediate access to necessary documents. However, there is a clear call for improvements, such as the ability to “Include the possibility of opening IO, SPL, etc. in .pdf format for consultation” (P3). This theme is supported by requests for direct links to full documents, suggesting that while “the list of relevant documents from which information is taken is excellent” (P4), the ability to delve deeper into full documents would significantly enhance the user experience.
Efficiency Users value the “greater speed in carrying out some small tasks” (P3). However, there are concerns about the system’s efficiency when it does not have the answer, leading to “wasting time looking for a solution to a problem in case it is not reported in the system’s history” (P3). Statements like “quick in responses” (P3) contrast with the need for questions to be “too specific to have a reliable answer” (P7), indicating tension between the desire for quick solutions and the system’s limitations.
Adoption Users highlight several factors affecting adopting the new system. It includes challenges such as “awareness and training of operators [might hinder adoption]” (P3) and the need for “acceptance by all employees” (P4), which indicates that the system’s success is contingent on widespread user buy-in. The generational divide is also noted: “That older operators use it [on what may hinder adoption]” (P7) suggests that demographic factors may influence the acceptance of new technology.
Safety Users express apprehension that “if the responses are not adequate, you risk safety” (P1), emphasizing the critical nature of reliable information in a factory setting. Moreover, the demand for updated and specific information underlines the importance of the system’s content being current and detailed to maintain operational safety standards, as stated by P9: “If it is updated and specific, it can help me”.
Traditional versus Novel There is a noticeable preference for established practices among some users. For instance, “It’s faster and easier to ask an expert colleague working near me rather than [the system]” (P8) captures the reliance on human expertise over the assistant system. This tension is further demonstrated by the sentiment that “Operators may benefit more from traditional information retrieval systems” (P9), suggesting a level of skepticism or comfort with the status quo that the new system needs to overcome.
5. LLM Benchmarking
@@
In our benchmarking experiment, we evaluated various commercial and open-source LLMs, including OpenAI’s ChatGPT (GPT-3.5 and GPT-4 from July 20th 2023), Guanaco 65B and 35B variants (Dettmers et al., 2023) based on Meta’s Llama (Large Language Model Meta AI) (Touvron et al., 2023a), Mixtral 8x7b (Jiang et al., 2024), Llama 2 (Touvron et al., 2023b), and one of its derivatives, StableBeluga2 (Mahan et al., 2023). This selection represents the state-of-the-art closed-sourced models (e.g., GPT-4) and open-source models (e.g., Llama 2). We included the (outdated) Guanaco models to demonstrate the improvements in the open-source sphere over the past year.
@@
We used a web UI for LLMs444https://github.com/oobabooga/text-generation-webui/tree/main—last accessed February 26, 2024 to load and test the Mixtral 8x7B, Guanaco models, and the StableBeluga2. The models were loaded on a pair of Nvidia A6000s with NVlink and a total Video Random Access Memory (VRAM) capacity of 96 GB. The 65B model was run in 8-bit mode to fit in the available VRAM. We used the llama-precise parameter preset and fixed zero seed for reproducibility. Llama 2 was evaluated using the demo on huggingface555https://huggingface.co/meta-llama/Llama-2-70b-chat-hf—last accessed February 26, 2024.
@@
To rigorously assess the models, we prepared 20 questions of varying complexity based on two types of context material: half from operating manuals and half from unstructured issue reports. The operating manuals included excerpts from actual machine manuals and standard operating procedures, while the informal issue reports were free-text descriptions of issues we had previously collected from operators. The model prompt was constructed using the above template (3.3). Ultimately, the difficulty of a question is a combination of the question’s complexity and the clarity of the source material. Simple questions include retrieving a single piece of information clearly stated in the context material, for example, ”At what temperature is relubrication necessary for the OKS 4220 grease?”. Conversely, difficult questions require more reasoning or comprise multiple parts, for example, ”What should I do if the central turntable is overloaded?” which has a nuanced answer dependent on several factors not clearly articulated in the context material.
In addition to measuring response length in words, every response is manually scored on factuality, completeness, and hallucinations as defined below:
Factuality: Responses align with the facts in the context material.
Completeness: Responses contain all the information relevant to the question in the context material.
Hallucinations: Response appears grammatically and semantically coherent but is not based on the context material.
@@
The following scoring protocol is applied: one is awarded for a completely factual, complete, or hallucinated response. In contrast, a score of 0.5 is awarded for a slightly nonfactual, incomplete, or hallucinated response (e.g., the response includes four out of the five correct steps). Otherwise, a score of zero is awarded. Therefore, wrong answers are penalized heavily. If the model responds by saying it cannot answer the question and does not make any attempt to do so, it is scored zero for factuality and completeness, but no score is given for hallucination. As such, the final score for hallucination is calculated as follows:
@@
As shown in Figure 3 and Table 1, GPT-4 outperforms other models regarding factuality, completeness, and lack of hallucinations but is closely followed by StableBeluga2 and GPT-3.5. The Guanaco models, based on Llama 1, perform significantly worse. The conciseness of the responses showed a similar pattern, except that StableBeluga2 produced the shortest answers (58 words), followed closely by Mixtral 8x7B (66 words) and GPT-4 (69 words).
@@
6. Discussion
6.1. GPT-4 is the best, but open-source models follow closely
@@
GPT-4 performs best across all measures but is closely followed by StableBeluga2, Mixtral 8x7B, and GPT-3.5. Compared to GPT-4, the cost per input token for GPT-3.5 is significantly lower666https://openai.com/pricing#language-models—last accessed February 26, 2024. However, the higher costs of GPT-4 are partially counteracted by its concise yet complete responses. If longer, more detailed responses were desired (e.g., for training purposes), the prompt could be adjusted. We observed that the less powerful models, such as GPT-3.5 and Llama 2, tended to be wordier and include additional details that were not directly requested. In contrast, GPT-4, StableBeluga2, and Mixtral 8x7B generated more concise responses.
@@
The latest generation of open-source models, such as Mixtral 8x7B and Llama 2 variants, such as StableBeluga2, demonstrates a clear jump forward relative to their predecessors based on Llama-1, which were more prone to hallucinations and exhibited poorer reasoning abilities over the context material. While open-source models like StableBeluga2 and Mixtral 8x7B do not score as high as GPT-4, they ensure better data security, privacy, and customization if hosted locally. This can be a crucial consideration for companies with sensitive data or unique needs.
6.2. The tool is beneficial but inferior to human experts
@@
Users appreciate the system’s functionality and see it as a tool for modernizing factory operations and speeding up operations. They are keen on improvements to be made for better user experience and utility, especially in the areas of content, feature enhancements, and user training. However, they express concerns about potential safety risks and the efficacy of information retrieval compared to consulting expert personnel. While these concerns are understandable, the tool was not designed to replace human-human interactions; instead, it can be used when no human experts are present or when they do not know or remember how to solve a specific issue. This would come into play during the night shift at the factory where we conducted the user study as a single operator operates a production line, leaving limited options for eliciting help from others.
6.3. Limitations and future work
@@
We used the same prompt for all LLMs; however, it is possible that some of the LLMs would perform better with a prompt template developed explicitly for it. However, we matched the LLMs’ hyperparameters (e.g., temperature) as closely as possible across all the tested models, except for Llama 2, as we did not have access to the presets as we did not host it locally. Our model benchmarking procedure involved 20 questions, and a singular coder assessed the responses. This introduces the potential for bias, and the limited number of questions may not cover the full spectrum of complexities in real-world scenarios. However, we varied query complexity and source material types to (partially) mitigate these shortcomings.
@@
The study’s design did not include a real-world evaluation involving end users operating the production line, as this was considered too risky for our industry partner. Such an environment might present unique challenges and considerations not addressed in this research, such as time pressure. However, by involving operators and managers and instructing them to pose several questions based on their actual work experience, we could still evaluate the system and collect valid feedback.
These limitations suggest directions for future research, for example, longitudinal studies where operators use the tool during production line operations and more comprehensive prompt and model customization. Longitudinal studies will be key to understanding the real-world impact on production performance, operator well-being, and cognitive abilities.
@@ 7. Conclusion
The results demonstrated GPT-4’s superior performance over other models regarding factuality, completeness, and minimal hallucinations. Interestingly, open-source models like StableBeluga2 and Mixtral 8x7B followed close behind. The user study highlighted the system’s user-friendliness, speed, and logical functionality. However, improvements in the user interface and content specificity were suggested, along with potential new features. Benefits included modernizing factory operations and speeding up specific tasks, though concerns about safety, efficiency, and inferiority to asking human experts were raised.
@@ Acknowledgements.
- Man (2014) 2014. The Manufacturing Knowledge Repository - Consolidating Knowledge to Enable Holistic Process Knowledge Management in Manufacturing:. In Proceedings of the 16th International Conference on Enterprise Information Systems (Lisbon, Portugal). SCITEPRESS - Science and and Technology Publications, 39–51. https://doi.org/10.5220/0004891200390051
@@
- Fantini et al. (2020) Paola Fantini, Marta Pinzone, and Marco Taisch. 2020. Placing the operator at the centre of Industry 4.0 design: Modelling and assessing human activities within cyber-physical systems. Computers & Industrial Engineering 139 (2020), 105058. https://doi.org/10.1016/j.cie.2018.01.025
- Feng et al. (2017) Shaw C. Feng, William Z. Bernstein, Jr. Thomas Hedberg, and Allison Barnard Feeney. 2017. Toward Knowledge Management for Smart Manufacturing. Journal of computing and information science in engineering 17 3 (2017). https://api.semanticscholar.org/CorpusID:4780425
@@
- Guest et al. (2011) Greg Guest, Kathleen M MacQueen, and Emily E Namey. 2011. Applied thematic analysis. sage publications.
@@
- Jiang et al. (2024) Albert Q. Jiang, Alexandre Sablayrolles, Antoine Roux, Arthur Mensch, Blanche Savary, Chris Bamford, Devendra Singh Chaplot, Diego de las Casas, Emma Bou Hanna, Florian Bressand, Gianna Lengyel, Guillaume Bour, Guillaume Lample, Lélio Renard Lavaud, Lucile Saulnier, Marie-Anne Lachaux, Pierre Stock, Sandeep Subramanian, Sophia Yang, Szymon Antoniak, Teven Le Scao, Théophile Gervet, Thibaut Lavril, Thomas Wang, Timothée Lacroix, and William El Sayed. 2024. Mixtral of Experts. arXiv:2401.04088 [cs.LG]
- Kernan Freire et al. (2023a) Samuel Kernan Freire, Mina Foosherian, Chaofan Wang, and Evangelos Niforatos. 2023a. Harnessing Large Language Models for Cognitive Assistants in Factories. In Proceedings of the 5th International Conference on Conversational User Interfaces (Eindhoven, Netherlands) (CUI ’23). Association for Computing Machinery, New York, NY, USA, Article 44, 6 pages. https://doi.org/10.1145/3571884.3604313
- Kernan Freire et al. (2023b) Samuel Kernan Freire, Chaofan Wang, Santiago Ruiz-Arenas, and Evangelos Niforatos. 2023b. Tacit Knowledge Elicitation for Shop-floor Workers with an Intelligent Assistant. In Extended Abstracts of the 2023 CHI Conference on Human Factors in Computing Systems. 1–7. https://doi.org/10.1145/3544549.3585755
@@
- Leoni et al. (2022) Luna Leoni, Marco Ardolino, Jamal El Baz, Ginetta Gueli, and Andrea Bacchetti. 2022. The Mediating Role of Knowledge Management Processes in the Effective Use of Artificial Intelligence in Manufacturing Firms. 42, 13 (2022), 411–437. https://doi.org/10.1108/IJOPM-05-2022-0282
@@
- Maddikunta et al. (2022a) Praveen Kumar Reddy Maddikunta, Quoc-Viet Pham, Prabadevi B, N Deepa, Kapal Dev, Thippa Reddy Gadekallu, Rukhsana Ruby, and Madhusanka Liyanage. 2022a. Industry 5.0: A survey on enabling technologies and potential applications. Journal of Industrial Information Integration 26 (2022), 100257. https://doi.org/10.1016/j.jii.2021.100257
- Maddikunta et al. (2022b) Praveen Kumar Reddy Maddikunta, Quoc-Viet Pham, B Prabadevi, Natarajan Deepa, Kapal Dev, Thippa Reddy Gadekallu, Rukhsana Ruby, and Madhusanka Liyanage. 2022b. Industry 5.0: A survey on enabling technologies and potential applications. Journal of Industrial Information Integration 26 (2022), 100257.
@@
- May et al. (2015) Gökan May, Marco Taisch, Andrea Bettoni, Omid Maghazei, Annarita Matarazzo, and Bojan Stahl. 2015. A New Human-centric Factory Model. 26 (2015), 103–108. https://doi.org/10.1016/j.procir.2014.07.112
@@
- Richter et al. (2019) Shahper Richter, Lena Waizenegger, Melanie Steinhueser, and Alexander Richter. 2019. Knowledge Management in the Dark: The Role of Shadow IT in Practices in Manufacturing. 15, 2 (2019), 1–19. https://doi.org/10.4018/IJKM.2019040101
@@
- Shneiderman (2022) Ben Shneiderman. 2022. Human-centered AI. Oxford University Press.
@@
- Wang et al. (2023a) Xingzhi Wang, Nabil Anwer, Yun Dai, and Ang Liu. 2023a. ChatGPT for design, manufacturing, and education. Procedia CIRP 119 (2023), 7–14. https://doi.org/10.1016/j.procir.2023.04.001 The 33rd CIRP Design Conference.
- Wang et al. (2023b) Zezhong Wang, Fan Yang, Pu Zhao, Lu Wang, Jue Zhang, Mohit Garg, Qingwei Lin, and Dongmei Zhang. 2023b. Empower Large Language Model to Perform Better on Industrial Domain-Specific Question Answering. ArXiv abs/2305.11541 (2023). https://api.semanticscholar.org/CorpusID:258822888
@@
- Wei et al. (2022a) Jason Wei, Yi Tay, Rishi Bommasani, Colin Raffel, Barret Zoph, Sebastian Borgeaud, Dani Yogatama, Maarten Bosma, Denny Zhou, Donald Metzler, Ed H. Chi, Tatsunori Hashimoto, Oriol Vinyals, Percy Liang, Jeff Dean, and William Fedus. 2022a. Emergent Abilities of Large Language Models. arXiv:2206.07682 [cs.CL]
- Wei et al. (2022b) Jason Wei, Xuezhi Wang, Dale Schuurmans, Maarten Bosma, brian ichter, Fei Xia, Ed Chi, Quoc V Le, and Denny Zhou. 2022b. Chain-of-Thought Prompting Elicits Reasoning in Large Language Models. In Advances in Neural Information Processing Systems, S. Koyejo, S. Mohamed, A. Agarwal, D. Belgrave, K. Cho, and A. Oh (Eds.), Vol. 35. Curran Associates, Inc., 24824–24837. https://proceedings.neurips.cc/paper_files/paper/2022/file/9d5609613524ecf4f15af0f7b31abca4-Paper-Conference.pdf
@@
- Wen et al. (2023) Cheng Wen, Xianghui Sun, Shuaijiang Zhao, Xiaoquan Fang, Liang Chen, and Wei Zou. 2023. ChatHome: Development and Evaluation of a Domain-Specific Language Model for Home Renovation. ArXiv abs/2307.15290 (2023). https://api.semanticscholar.org/CorpusID:260315830
@@
- Xie et al. (2023a) Qianqian Xie, Weiguang Han, Xiao Zhang, Yanzhao Lai, Min Peng, Alejandro Lopez-Lira, and Jimin Huang. 2023a. PIXIU: A Large Language Model, Instruction Data and Evaluation Benchmark for Finance. (6 2023). https://arxiv.org/abs/2306.05443v1
- Xie et al. (2023b) Tong Xie, Yuwei Wan, Wei Huang, Zhenyu Yin, Yixuan Liu, Shaozhou Wang, Qingyuan Linghu, Chunyu Kit, Clara Grazian, Wenjie Zhang, Imran Razzak, and Bram Hoex. 2023b. DARWIN Series: Domain Specific Large Language Models for Natural Science. ArXiv abs/2308.13565 (2023). https://api.semanticscholar.org/CorpusID:264960172
@@
- Xu et al. (2021a) Xun Xu, Yuqian Lu, Birgit Vogel-Heuser, and Lihui Wang. 2021a. Industry 4.0 and Industry 5.0—Inception, conception and perception. Journal of Manufacturing Systems 61 (2021), 530–535.
- Xu et al. (2021b) Xun Xu, Yuqian Lu, Birgit Vogel-Heuser, and Lihui Wang. 2021b. Industry 4.0 and Industry 5.0—Inception, conception and perception. Journal of Manufacturing Systems 61 (2021), 530–535. https://doi.org/10.1016/j.jmsy.2021.10.006
- Zhang et al. (2023a) Jianzhang Zhang, Yiyang Chen, Nan Niu, and Chuang Liu. 2023a. A Preliminary Evaluation of ChatGPT in Requirements Information Retrieval. (4 2023). https://arxiv.org/abs/2304.12562v1
- Zhang et al. (2023b) Wenxuan Zhang, Hongzhi Liu, Yingpeng Du, Chen Zhu, Yang Song, Hengshu Zhu, and Zhonghai Wu. 2023b. Bridging the Information Gap Between Domain-Specific Model and General LLM for Personalized Recommendation. ArXiv abs/2311.03778 (2023). https://api.semanticscholar.org/CorpusID:265043714
@@
- Zuccon et al. (2023) Guido Zuccon, Bevan Koopman, and Razia Shaik. 2023. ChatGPT Hallucinates when Attributing Answers. In Proceedings of the Annual International ACM SIGIR Conference on Research and Development in Information Retrieval in the Asia Pacific Region (¡conf-loc¿, ¡city¿Beijing¡/city¿, ¡country¿China¡/country¿, ¡/conf-loc¿) (SIGIR-AP ’23). Association for Computing Machinery, New York, NY, USA, 46–51. https://doi.org/10.1145/3624918.3625329